OpenStudy (anonymous):

hey so i dont really understand f of x. i mean i kinda do but then i get a problem like this Let f(x) = x + 8 and g(x) = x2 - 6x - 7. Find f(g(2)). and im so confused! please help

OpenStudy (anonymous):

do you have multiple choice answers? if so please show them :)

hartnn (hartnn):

to get g(2) you just need to plug in x=2 in g(x) what would you get as g(2) ? try...

OpenStudy (anonymous):

im in algebra 2 and im good at math but this is the one thing i never understood.

OpenStudy (anonymous):

plus im a little rusty

OpenStudy (anonymous):

((2)2-6(2)-7)+8

hartnn (hartnn):

example if \(f(x) = x^2+ 4x \\ f(\heartsuit) = \heartsuit^2 +4 \heartsuit \)

hartnn (hartnn):

\(g(x) = x^2-6x-7 \\ g(2) = ... ?\)

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{ f(x) = x + 8}\) \(\normalsize\color{blue}{g(x) = x^2 - 6x - 7}\) \(\normalsize\color{blue}{f(g(x)) = (x^2 - 6x - 7) + 8}\) \(\normalsize\color{blue}{f(g(x)) = x^2 - 6x - 7 + 8}\) \(\normalsize\color{blue}{f(g(x)) = x^2 - 6x + 1}\) \(\normalsize\color{blue}{f(g(2)) = (2)^2 - 6(2) + 1}\) \(\normalsize\color{blue}{f(g(2)) =4 - 12+ 1}\) \(\normalsize\color{blue}{f(g(2)) =??}\) you tell me

OpenStudy (anonymous):

Plug in 2 in g(x) and then plug g(x) in the x if f(x) ..

OpenStudy (nirmalnema):

y=f(x) is y as a function of x f(x) = x + 8 and g(x) = x2 - 6x - 7. then f(g(2)) f(g(x)).= x^2 - 6x - 7. + 8 put g(x) in place of x f(g(2))=2^2 - 6*2 - 7. + 8 4+8-7-12 =-7

OpenStudy (anonymous):

Of*

OpenStudy (solomonzelman):

I would normally imagine something like g(f(x)) g of f, not f of g (in this case) ....

OpenStudy (anonymous):

i have processing problems so it may take me a second

OpenStudy (solomonzelman):

sure

OpenStudy (anonymous):

@SolomonZelman i got -7. i understand it now!!!! thank you and everyone else for there help

OpenStudy (solomonzelman):

-7 is right! ... Anytime!